# What Makes Math Word Problems Challenging for LLMs?

### v1

@@ Abstract
This paper investigates the question of what makes math word problems (MWPs) challenging for large language models (LLMs). We conduct an in-depth analysis of the key linguistic and mathematical characteristics of MWPs. In addition, we train feature-based classifiers to better understand the impact of each feature on the overall difficulty of MWPs for prominent LLMs and investigate whether this helps predict how well LLMs fare against specific categories of MWPs.
@@ 1 Introduction
A recent study by Almoubayyed et al. (2023) demonstrates a strong connection between reading skills and math outcomes in students.
We hypothesize that LLMs’ ability to solve MWPs correctly may similarly rely on: (1) the linguistic complexity of the questions; (2) the conceptual complexity of the tasks (e.g., the number of steps and types of math operations involved); and (3) the amount of real-world knowledge required to solve the tasks. Supporting this intuition, our preliminary analysis of the GSM8K dataset Cobbe et al. (2021) suggests that relatively short questions with a small number of described entities, a few calculation steps and a limited range of operators involved in the solution (e.g., Mark is 7 years older than Amy, who is 15. How old will Mark be in 5 years?) are typically answered correctly by a range of LLMs. At the same time, long questions requiring real-world knowledge (e.g., how many cents there are in a dollar) and extended natural language understanding (NLU) (e.g., interpretation of a lower price) pose challenges for LLMs (see Figure 1).
In this paper, we formulate and investigate two research questions: (1) Which characteristics of the input math word question make it complex for an LLM? and (2) Based on these characteristics, can we predict whether a particular LLM will be able to solve specific input MWPs correctly?
@@ 2 Methodology
We use the GSM8K dataset  Cobbe et al. (2021), divided into $7,473$ training and $1,319$ test instances, because of the high quality of human-generated MWPs. This dataset contains a diverse set of problems in English with minimal amount of recurring templates. Further, the difficulty level of the problems is tailored for LLMs, allowing for a wide variation in correctness across models and question types, which is ideal for our feature-based analysis.
@@ 3.2 Success Rate Prediction
For high confidence samples, we use the training and test subset from GSM8K where the sampled success rate is either $1.0$ (always correct) or $0.0$ (never correct). The distribution of the LLM-specific splits are detailed in Table 2.
@@ 4 Results
Train Always 1102 (30.22%) 2438 (61.36%) 733 (24.06%) 5162 (94.7%) 214 (53.37%)
Never 2545 (69.78%) 1535 (38.64%) 2313 (75.94%) 289 (5.3%) 187 (46.63%)
@@
Test Always 188 (28.14%) 427 (60.06%) 111 (21.51%) 528 (71.64%) 41 (30.37%)
Never 480 (71.86%) 284 (39.94%) 405 (78.49%) 209 (28.36%) 94 (69.63%)
@@
Table 2: Class-wise distribution and classification results for different LLMs. "Intersection" refers to questions that were always or never solved correctly by all or any LLM respectively. All classification results are mean values across 5 runs with varying initialization seeds.
@@ 4.1 Success Rate Distribution
We report the mean success rates for each LLM on GSM8K’s test set in Table 1.222Our results generally align with those reported previously for these models. We observe that Llama2 13B and 70B follow the expected order of scores along their respective parameter counts. Mistral-7B scores similar to the 13B Llama2 models, and the additional fine-tuning allows MetaMath-13B to outperform the other models (including the 70B Llama2). Figures 1(a) and 1(b) respectively capture the number of questions always and never answered correctly by each LLM. Overall, MetaMath-13B has the lowest number of incorrectly and the highest number of correctly answered questions across the tested LLMs.
@@
Figure 2: Number of questions from the GSM8K-test (a) always and (b) never answered correctly by each LLM. The rows in either figures correspond to individual LLMs, with the counts on the right denoting the total number of questions always (or never) answered correctly by each LLM. The counts at the bottom denote the number of questions in each subset of LLMs.
@@ 4.2 Classification Results
At the same time, we also note that, due to significant class imbalance, this task is not easy for the classifiers, with the best accuracy scores across LLM splits being in the range of $71.7\%-81.4\%$. The small number of questions always or never solved correctly by any LLM speaks to the varying capabilities (and potential points of brittleness) of the models. We include additional analysis of the results in Appendix D.
@@ 4.2.1 Feature Importance
The statistical classifiers used in our experiments allow us to estimate the importance of each feature and its contribution to the classification performance. We report the top $10$ features with the highest aggregate ranks across LLM data splits and classifiers in Table 3. We use mean rank here as a proxy for relative importance between features and the respective standard deviations indicate how spread-out this importance is across classifiers and queried LLMs. We observe that a greater number (Gx_op_unique_count) and diversity (Gx_op_diversity) in math operations, and the use of infrequent numerical tokens in the question and solution body (Qx_ & Gx_mean_numerical_word_rank) impact the success rate. The list also contains linguistic features based on the phrasing of the questions: longer questions with a high number of noun phrases (Qx_np_count), mean syntactic depth (Qx_constituency_tree_depth), and readability grade (Qx_flesch_kinkaid_grade) are also difficult for LLMs to solve. Additionally, the need for extraneous information (Gx_world_knowledge) such as conversion units for time, distance, or weight, can make a question challenging. We also report value thresholds at which each feature affects the success rate significantly: see the results of the Student’s t-test and p-values in Table 7 in Appendix D.
@@ 4.2.2 Ablation Studies
To further measure the impact of each feature type, we report classification scores along different feature-type subsets in Figure 3. We note that the feature set with all types (L+M+W) is not the most optimal for classification. For instance, the questions answered by Llama2-13B are best classified using only mathematical features (M). The best performing classifiers for Llama2-7B, MetaMath-13B, and the intersection set either solely use linguistic features (L), or both linguistic and math features (L+M), whereas the world knowledge & NLU feature set if sufficient for Mistral-7B.
Figure 3: Results of the ablation studies across feature types (L - Linguistic, M - Mathematical, W - World Knowledge & NLU). Each bar represents the mean macro-F1 score over all three classifier models.
@@ 4.2.3 Impact of Linguistic Features
In order to better gauge the impact of linguistic features on the success rate, we cluster questions by mathematical features. We fit a KMeans clustering model on all math features for each question in the GSM8K train-set with a target cluster count of $100$. This helps group together questions from the data, wherein the math features hardly vary within each subset (or cluster) of questions. Thus, variations in success rate across the questions within a cluster can be more clearly attributed to other, i.e., linguistic types of features. We report some notable Spearman correlation values between the linguistic feature values within a cluster and the corresponding success rates in Table 4. The strong and significant feature-wise negative correlations suggest that for a relatively fixed set of math features, questions with greater length, nesting, lexical rank, and reading grade become more challenging for LLMs to solve. Note that this form of analysis on feature-based minimal pairs is extractive in nature and may, to a certain extent, be restricted to the question types in the GSM8K dataset. For a more exhaustive analysis for each feature, generative approaches to furnish question paraphrases with the desired set of linguistic features need to be employed.
@@ 5 Conclusions
This work aims to identify what aspects of MWPs make them difficult for LLMs to solve. To this end, we extract key features (spanning linguistic, mathematical, and real-world knowledge & NLU based aspects) to predict whether MWPs from GSM8K can be reliably solved by several LLMs. [^3] Our data and analysis are publicly available at github.com/kvadityasrivatsa/analyzing-llms-for-mwps We find that questions with a high number and diversity of math operations using infrequent numerical tokens are particularly challenging to solve. In addition, we also show that lengthy questions with low readability scores and those requiring real-world knowledge are also seldom solved correctly. Our future work will rely on these findings to make informed modifications to questions in order to study the impact on LLMs’ reasoning and MWP solving abilities. Figure 4 provides an example of an informed modification, which leads to improved LLM performance.
@@
We consider it important to test and report results with a diverse set of open-source LLMs, which motivated the selection of the specific models included in this study. At the same time, we do not claim this study to be comprehensive with respect to the range of LLMs tested and, in future work, we plan to include more LLMs in this research.
@@
In this work, we have focused on a single MWP dataset (GSM8K) due to its unique properties, namely high quality of the questions, high diversity of the tasks (including linguistic diversity of the questions), and moderate difficulty of the math problems covered Cobbe et al. (2021). At the same time, we recognize that the results that we report in this work may be limited in certain ways to the dataset that we report them on. Our future work will apply this approach to other available MWP datasets Kim et al. (2023); Wang and Lu (2023); Patel et al. (2021); Miao et al. (2020) to verify the consistency of the findings.
@@
Finally, whilst we have identified aspects of the MWPs that make them challenging for LLMs to solve, we admittedly presented only one example (see Figure 4) where acting upon one of the identified aspects improves the output of an LLM. While a thorough investigation of the impact of such informed modifications is outside the scope of the current paper, such experiments will follow in future work to demonstrate practical usefulness of the identified MWP aspects.
@@ Appendix A Feature Details
Below we describe the features used in our study and the way they were extracted.
@@
Table 5: Details of formulation and distribution (across GSM8K) for all features included in the feature set. Each feature is of type: Linguistic (L), Mathematical (M), or World Knowledge and NLU (W); and is sourced either from the question body (Q), gold solution body (G), or both (B).
@@
  • Qx_token_length: The number of tokens in the tokenized version of the question body. We apply each LLM’s respective tokenizer from HuggingFace to collect this feature.
@@
  • Qx_prp_count: Number of prepositions found in the question body. We use the part-of-speech tags generated as part of the parse by Stanza.
  • Qx_coref_count: Number of pronomial or nominal instances of coreference in the question body. We use Stanford’s CorefAnnotator to extract this feature.
@@
  • Qx_arg_count: The number of distinct numerical quantities (e.g., “3.5 hours later” or “100 boxes”) in the question body. We use a Regexp pattern to detect whole numbers, decimal point values, quantities preceded by a negative sign or dollar (and other currency) signs.
  • Qx_word_arg_count: The number of quatities mentioned in word-form (“three times” or “half as much”) in the question body. We use a vocabulary of frequently used word-form tokens and accommodate compound expressions (e.g., “twenty-two”).
@@
  • Gx_arg_count: The number of distinct numerical quantities present as plain text or in the left-hand side of equations in the gold solution. We use the same Regexp pattern used for Qx_arg_count.
@@
  • Gx_op_diversity: Ratio of the number of unique math operators used to the total number of operators present in the gold solution.
  • Gx_mean_numerical_word_rank The mean vocabulary rank of the numerical tokens used on the left-hand-side of equations in the gold solution. Extracted the same way as Qx_mean_numerical_word_rank.
  • Bx_parameter_usage The ratio of the number of distinct arguments used in the gold solution to that in the question body. A value lower than 1 indicates that one or more arguments provided in the question were not required to solve the MWP (potentially acting as distractors).
(3) World Knowledge and NLU (W) features include:
@@
Figure 5: Spearman correlation matrix between features. All correlational values are marked with ‘*’, ‘**’, and ‘***’ if their corresponding p-values are less than $0.05$, $0.01$, and $0.001$ respectively.
@@ B.2 LLM Details
Llama2-13B meta-llama/Llama-2-13b-chat-hf 28.7
Llama2-70B meta-llama/Llama-2-70b-chat-hf 56.8
@@
MetaMath-13B meta-math/MetaMath-13B-V1.0 72.3
Table 6: List of Huggingface model variants and their respective reported pass@1 (single run) accuracy on the GSM8K test set by the OpenLLM leader board.
@@ B.3 Implementation and Compute Resources Used
We use VLLM to load and query models. Models of parameter size 7B and 13B were queried with a single NVIDIA A100 GPU. Llama2-70B was loaded and queried using 4 A100 GPUs. Each query was set to a temperature of $0.8$, and a maximum token length of $2000$. Each question was queried $5$ times by each LLM, with a varying seed. Querying the entire GSM8K dataset ($8,793$ questions) took approximately 1 hour for each LLM.
@@
Table 7: Feature-wise thresholds which reflect the greatest difference in the corresponding mean success rate. For each feature, the optimal threshold creates two sets of questions on either side, wherein the difference in the corresponding mean success rates of the two sets is greatest. We perform Student’s T-Tests on both sets to determine if this difference is significant and report the corresponding T values. All results reported in the table have an absolute T-value >5 and a p-value <0.0001.

### v2

@@ Abstract
This paper investigates the question of what makes math word problems (MWPs) in English challenging for large language models (LLMs). We conduct an in-depth analysis of the key linguistic and mathematical characteristics of MWPs. In addition, we train feature-based classifiers to better understand the impact of each feature on the overall difficulty of MWPs for prominent LLMs and investigate whether this helps predict how well LLMs fare against specific categories of MWPs. [^1] Our code, data, and analysis are publicly available at github.com/kvadityasrivatsa/analyzing-llms-for-mwps
@@ 1 Introduction
A recent study by Almoubayyed et al. (2023) demonstrates a strong connection between reading skills and math outcomes in students. We hypothesize that LLMs’ ability to solve MWPs correctly may similarly rely on: (1) the linguistic complexity of the questions; (2) the conceptual complexity of the tasks (e.g., the number of steps and types of math operations involved); and (3) the amount of real-world knowledge required to solve the tasks. Supporting this intuition, our preliminary analysis of the GSM8K dataset Cobbe et al. (2021) suggests that relatively short questions with a small number of described entities, a few calculation steps and a limited range of operators involved in the solution (e.g., Mark is 7 years older than Amy, who is 15. How old will Mark be in 5 years?) are typically answered correctly by a range of LLMs. At the same time, long questions requiring real-world knowledge (e.g., how many cents there are in a dollar) and extended natural language understanding (NLU) (e.g., interpretation of a lower price) pose challenges for LLMs (see Figure 1).
In this paper, we formulate and investigate two research questions: (1) Which characteristics of the input math word questions make them complex for an LLM? and (2) Based on these characteristics, can we predict whether a particular LLM will be able to solve specific input MWPs correctly?
@@ 2 Methodology
We use the GSM8K dataset  Cobbe et al. (2021), divided into $7,473$ training and $1,319$ test instances, because of the high quality of human-generated MWPs. This dataset contains a diverse set of problems in English with minimal amount of recurring templates. Furthermore, the difficulty level of the problems is tailored for LLMs, allowing for a wide variation in correctness across models and question types, which is ideal for our feature-based analysis.
@@ 3.2 Success Rate Prediction
For high confidence samples, we use the training and test subset from GSM8K where the sampled success rate is either $1.0$ (always correct) or $0.0$ (never correct). The distribution of the LLM-specific splits is detailed in Table 2.
@@ 4 Results
Train Always 1102 (30.22%) 2438 (61.36%) 733 (24.06%) 5162 (94.7%) 205 (53.38%)
Never 2545 (69.78%) 1535 (38.64%) 2313 (75.94%) 289 (5.3%) 179 (46.61%)
@@
Test Always 188 (28.14%) 427 (60.06%) 111 (21.51%) 528 (71.64%) 31 (24.41%)
Never 480 (71.86%) 284 (39.94%) 405 (78.49%) 209 (28.36%) 96 (75.59%)
@@
RoBERTa-base 0.816 0.771 0.756 0.738 0.838 0.743 0.701 0.415 0.811 0.781
Table 2: Class-wise distribution and classification results for different LLMs. "Intersection" refers to questions always or never solved correctly by all or any LLM, respectively. All classification results are mean values across 5 runs with varying initialization seeds. The best results for feature-based classifiers are highlighted in bold.
@@ 4.1 Success Rate Distribution
We report the mean success rates for each LLM on GSM8K’s test set in Table 1.333Our results generally align with those reported previously for these models. We observe that Llama2 13B and 70B follow the expected order of scores along their respective parameter counts. Mistral-7B scores similar to the 13B Llama2 model, and the additional fine-tuning allows MetaMath-13B to outperform the other models (including the 70B Llama2). Figures 2(a) and 2(b) respectively capture the number of questions always and never answered correctly by each LLM. Overall, MetaMath-13B has the lowest number of incorrectly and the highest number of correctly answered questions across the tested LLMs.
@@
Figure 2: Number of questions from the GSM8K-test (a) always and (b) never answered correctly by each LLM. The rows in each figure correspond to individual LLMs, with the counts on the right denoting the total number of questions always (or never) answered correctly by each LLM. The counts at the bottom denote the number of questions in each subset of LLMs.
@@ 4.2 Classification Results
At the same time, we also note that, due to significant class imbalance, this task is not easy for the classifiers, with the best accuracy scores across LLM splits being in the range of $71.7\%-81.4\%$. The small number of questions always or never solved correctly by any LLM speaks to the models’ varying capabilities (and potential points of brittleness). We include additional analysis of the results in Appendix D.
For comparison, we also report the classification results for a fine-tuned RoBERTa-base model Liu et al. (2019) for the same training and evaluation sets (tuned on the question and gold solution as input text; see Appendix C for more details) in Table 2. We note that the Transformer base classifier scores on a par or a few points above the best statistical classifier, i.e., Random Forest, suggesting that the proposed feature-based classifiers are not far behind token-level contextual models for this task.
@@ 4.2.1 Feature Importance
The statistical classifiers used in our experiments allow us to estimate the importance of each feature and its contribution to the classification performance. We report the top $10$ features with the highest aggregate ranks across LLM data splits and classifiers in Table 3. We use mean rank here as a proxy for relative importance across features, and the respective standard deviations indicate how spread out this importance is across classifiers and queried LLMs. We observe that a greater number (Gx_op_unique_count) and diversity (Gx_op_diversity) in math operations, and the use of infrequent numerical tokens in the question and solution body (Qx_ & Gx_mean_numerical_word_rank) impact the success rate. The list also contains linguistic features based on the phrasing of the questions: longer questions with a high number of noun phrases (Qx_np_count), mean syntactic depth (Qx_constituency_tree_depth), and readability grade (Qx_flesch_kinkaid_grade) are also difficult for LLMs to solve. Additionally, the need for extraneous information (Gx_world_knowledge), such as conversion units for time, distance, or weight, can make a question challenging. We also report value thresholds at which each feature affects the success rate significantly: see the results of the Student’s t-test and p-values in Table 7 in Appendix D.
@@ 4.2.2 Ablation Studies
To further measure the impact of each feature type, we report classification scores along different feature-type subsets in Figure 3. We note that the feature set with all types (L+M+W) is not optimal for classification. For instance, the questions answered by Llama2-13B are best classified using only mathematical features (M). The best-performing classifiers for Llama2-7B, MetaMath-13B, and the intersection set either solely use linguistic features (L) or both linguistic and math features (L+M), whereas the world knowledge & NLU feature set if sufficient for Mistral-7B.
Figure 3: Results of the ablation studies across feature types (L – Linguistic, M – Mathematical, W – World Knowledge & NLU). Each bar represents the mean macro-F1 score over all three classifier models.
@@ 4.2.3 Impact of Linguistic Features
In order to better gauge the impact of linguistic features on the success rate, we cluster questions by mathematical features. We fit a KMeans clustering model [^4] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html on all math features for each question in the GSM8K training set with a target cluster count of $100$. This helps group together questions from the data, wherein the math features hardly vary within each question subset (or cluster). Thus, variations in success rate across the questions within a cluster can be more clearly attributed to other, i.e., linguistic types of features. We report some notable Spearman correlation values between the linguistic feature values within a cluster and the corresponding success rates in Table 4. The strong and significant feature-wise negative correlations suggest that for a relatively fixed set of math features, questions with greater length, nesting, lexical rank, and reading grade become more challenging for LLMs to solve. Note that this form of analysis on feature-based minimal pairs is extractive in nature and may, to a certain extent, be restricted to the question types in the GSM8K dataset. For a more exhaustive analysis for each feature, generative approaches to furnish question paraphrases with the desired set of linguistic features need to be employed.
@@ 5 Conclusions
This work aims to identify what aspects of MWPs make them difficult for LLMs to solve. To this end, we extract key features (spanning linguistic, mathematical, and real-world knowledge & NLU-based aspects) to predict whether several LLMs can reliably solve MWPs from GSM8K. We find that questions with a high number and diversity of math operations using infrequent numerical tokens are particularly challenging to solve. In addition, we show that lengthy questions with low readability scores and those requiring real-world knowledge are also seldom solved correctly. Our future work will rely on these findings to make informed modifications to questions in order to study the impact on LLMs’ reasoning and MWP-solving abilities. Figure 4 provides an example of an informed modification, which leads to improved LLM performance.
@@
We consider it important to test and report results with a diverse set of open-source LLMs, which motivated the selection of the specific models included in this study. At the same time, we do not claim this study to be comprehensive with respect to the range of LLMs tested and in future work, we plan to include more LLMs in this research.
@@
In this work, we have focused on a single MWP dataset (GSM8K) due to its unique properties, namely the high quality of the questions, high diversity of the tasks (including linguistic diversity of the questions), and moderate difficulty of the math problems covered Cobbe et al. (2021). At the same time, we recognize that the results we report in this work may be limited in certain ways to the dataset on which we report them. Our future work will apply this approach to other available MWP datasets Kim et al. (2023); Wang and Lu (2023); Patel et al. (2021); Miao et al. (2020) to verify the consistency of the findings.
@@
Finally, whilst we have identified aspects of the MWPs that make them challenging for LLMs to solve, we admittedly presented only one example (see Figure 4) where acting upon one of the identified aspects improves the output of an LLM. While a thorough investigation of the impact of such informed modifications is outside the scope of the current paper, such experiments will follow in future work to demonstrate the practical usefulness of the identified MWP aspects.
@@
Acknowledgments
We are grateful to the Campus Super Computing Center at MBZUAI for supporting this work. We also thank the anonymous reviewers for their valuable feedback.
@@ Appendix A Feature Details
Below, we describe the features used in our study and how they were extracted.
@@
Table 5: Details of formulation and distribution (across GSM8K) for all features included in the feature set. Each feature is of type: Linguistic (L), Mathematical (M), or World Knowledge and NLU (W) and is sourced either from the question body (Q), gold solution body (G), or both (B).
@@
  • Qx_token_length: The number of tokens in the tokenized version of the question body. We apply each LLM’s respective tokenizer from HuggingFace to extract this feature.
@@
  • Qx_prp_count: Number of prepositions in the question body. We use the part-of-speech tags generated as part of the parse by Stanza.
  • Qx_coref_count: Number of pronominal or nominal instances of coreference in the question body. We use Stanford’s CorefAnnotator to extract this feature.
@@
  • Qx_arg_count: The number of distinct numerical quantities (e.g., “3.5 hours later” or “100 boxes”) in the question body. We use a Regexp pattern to detect whole numbers, decimal point values, and quantities preceded by a negative sign or dollar (and other currency) signs.
  • Qx_word_arg_count: The number of quantities mentioned in word-form (“three times” or “half as much”) in the question body. We use a vocabulary of frequently used word-form tokens and accommodate compound expressions (e.g., “twenty-two”).
@@
  • Gx_arg_count: The number of distinct numerical quantities present as plain text or on the left-hand side of equations in the gold solution. We use the same Regexp pattern used for Qx_arg_count.
@@
  • Gx_op_diversity: Ratio of the number of unique math operators used to the total number of operators in the gold solution. For instance, a question with the consolidated math solution expression "$(2\times 12)\times 3=72$" contains two arithmetic operations in total but only one unique operation type, i.e., ‘$\times$,’ Gx_op_diversity$=1/2=0.5$.
  • Gx_mean_numerical_word_rank: The mean vocabulary rank of the numerical tokens used on the left-hand side of equations in the gold solution. Extracted the same way as Qx_mean_numerical_word_rank.
  • Bx_parameter_usage: The ratio of distinct arguments used in the gold solution to that in the question body. A value lower than 1 indicates that one or more arguments provided in the question were not required to solve the MWP (potentially acting as distractors).
(3) World knowledge and NLU features (W) include:
@@
Figure 5: Spearman correlation matrix between features. All correlation values are marked with ‘*’, ‘**’, and ‘***’ if their corresponding p-values are less than $0.05$, $0.01$, and $0.001$ respectively.
@@ B.2 LLM Details
Llama2-13B meta-llama/Llama-2-13b-chat-hf 28.70
Llama2-70B meta-llama/Llama-2-70b-chat-hf 56.80
@@
MetaMath-13B meta-math/MetaMath-13B-V1.0 72.30
Table 6: List of HuggingFace model variants and their respective reported pass@1 (single run) accuracies on the GSM8K test set from the OpenLLM leaderboard.
@@ B.3 Implementation and Compute Resources Used
We use vLLM to load and query models. Models of parameter sizes 7B and 13B were queried with a single NVIDIA A100 GPU. Llama2-70B was loaded and queried using 4 A100 GPUs. Each query was set to a temperature of $0.8$, and a maximum token length of $2000$. Each question was queried $5$ times by each LLM, with a varying seed. Querying the entire GSM8K dataset ($8,793$ questions) took approximately 1 hour for each LLM.
@@
Table 7: Feature-wise thresholds which reflect the greatest difference in the corresponding mean success rate. For each feature, the optimal threshold creates two sets of questions on either side, wherein the difference in the corresponding mean success rates of the two sets is the greatest. We perform Student’s $t$-tests on both sets to determine if this difference is significant and report the corresponding $t$ values. All results reported in the table have an absolute $t$-value >5 and a $p$-value <0.0001.